OpenStudy (anonymous):

write an arithmetic sequence that has 3 arithmetic means between -1/4 and -19/12

OpenStudy (anonymous):

@jishan can u help?

OpenStudy (anonymous):

@ChmE can u help please?

OpenStudy (anonymous):

@surdawi can u help??

OpenStudy (anonymous):

@supercrazy92 do you know this one? if so can u help?

OpenStudy (klimenkov):

Solve this: \(-\frac{19}{12}+d+d+d=-\frac14\)

OpenStudy (anonymous):

so the difference is -1?

OpenStudy (anonymous):

how do i find the difference? i used the formula but i'm not getting the right answer

OpenStudy (klimenkov):

Do you have an answer?

OpenStudy (anonymous):

no i dont, everytime i keep doing it i end up messing up

OpenStudy (anonymous):

here is what i got so far 19/12 -(-1/4) = -19/12 + 1/4 = -19/12 + 3/12 = -16/12 = -4. Since the difference is -4, divide by 4, = -1.

OpenStudy (anonymous):

are u there?

OpenStudy (klimenkov):

\(-\frac{16}{12}\ne -4\)

OpenStudy (anonymous):

so the difference is -4 right?

OpenStudy (anonymous):

???

OpenStudy (klimenkov):

\(-\frac54, -\frac{11}{12},-\frac7{12}\). Difference is \(\frac4{12}\).

OpenStudy (anonymous):

hmm how did u get it tho? can u explain?

OpenStudy (anonymous):

so the 3 terms will be -5/4, -11/12, and -7/12 ?

OpenStudy (klimenkov):

Yes.